OpenStudy (elisaneedshelp):

A rectangular garden 6 meters wide has a diagonal measuring 10 meters. Find the perimeter of the garden

OpenStudy (kc_kennylau):



OpenStudy (kc_kennylau):

Use the Pyth Thm perhaps?

OpenStudy (ipwnbunnies):

Right, use Pyth Thm to find the length of the rectangle. I trust you can find the perimeter knowing the length and width of a rectangle.

OpenStudy (elisaneedshelp):

does the diagonal count as width?....

OpenStudy (anonymous):

yeah right through pythagorian theorem u can solve it so you can get 8m and the perimeter is 28m

OpenStudy (ipwnbunnies):

No, they already gave you the width of 6. The diagonal is a diagonal.

OpenStudy (anonymous):

8 is the length

OpenStudy (elisaneedshelp):

How did you figure out the lenth? @heidi_xingxing

OpenStudy (elisaneedshelp):

@iPwnBunnies how did she figure out the lenght?

OpenStudy (ipwnbunnies):

The pythagorean theorem, manipulating it. a^2 + b^2 = c^2. We're trying to find a or b, however you're looking at it. a = sqrt(c^2 - b^2). a will be the length of the rectangle.

OpenStudy (elisaneedshelp):

but i dont understand how to do it in this problem

OpenStudy (ipwnbunnies):

They gave us this diagonal so we can find the length of the side.When the rectangle is split by the diagonal, it forms 2 right triangles. Coincidently, one side of the triangle is the width, and the other side of the triangle is the length, and they gave us the hypotenuse.

OpenStudy (anonymous):

\[\sqrt{10^{2}-6^{2}}\]

OpenStudy (anonymous):

given the width 6 and the length 8 from the pyth you can now solve the perimeter and P=28m

OpenStudy (kc_kennylau):